OpenStudy (anonymous):

What is the slope intercept form that passes through the given point and is perpendicular (3,2), y=x+5

OpenStudy (zale101):

find the negative reciprocal of the slope

OpenStudy (anonymous):

Can someone walk though it with me

OpenStudy (zale101):

and once you do, plug in the negative reciprocal of the equation along with the given point (3,2) into the point slope formula

OpenStudy (zale101):

so what is the negative reciprocal of the slope?

OpenStudy (anonymous):

-5? Y=x+5

OpenStudy (zale101):

y=x+5 1x is the reciprocal, if x is left alone that means there's an invisible 1 and it is the slope of the equation

OpenStudy (anonymous):

Ok then wut do I do?

OpenStudy (zale101):

the negative reciprocal of 1x the slope is -1 reciprocal means Interchanged

OpenStudy (zale101):

now what's next?

OpenStudy (zale101):

i mentioned the next step earlier :)

OpenStudy (anonymous):

But wut do I do with the (3,2) part

OpenStudy (zale101):

you'll plug in the negative reciprocal of 1x which is -1 and the given point (3,2) into the point slope formula. Do you still remember the point slope formula? oh and by the way, perpendicular is the reverse of the given slope but with parallel lines, the slope is the same it is not changed

OpenStudy (anonymous):

Is the answer y=3x -1

OpenStudy (zale101):

no

OpenStudy (zale101):

what is the point slope formula?

OpenStudy (anonymous):

?

OpenStudy (zale101):

y-y1=m(x-x1) m is the slope which is -1 x1,y1 is the given point (3,2)

OpenStudy (zale101):

can you plug those in the formula

OpenStudy (anonymous):

Y-2=-1(x-3)

OpenStudy (zale101):

yes, now distribute -1 into x-3 to isolate the parenthesis

OpenStudy (anonymous):

Y= -1x-3 ?

OpenStudy (zale101):

no, -1*x = -1x and -1*-3= +3 combine them -1x+3

OpenStudy (zale101):

clear?

OpenStudy (zale101):

so y-2=-1x+3

OpenStudy (zale101):

Now, what's next?

OpenStudy (zale101):

what should you do to isolate y?

OpenStudy (anonymous):

Is the answer y= -1x+1